OpenStudy (anonymous):

Differential Equations home work:

OpenStudy (anonymous):

So, there is no way for you to solve this.

OpenStudy (anonymous):

Find solution to the IVP and give in standard conic form \[\left( y ^{2}+16 \right)dx + \left( 3y-xy \right)dy = 0 \]

OpenStudy (anonymous):

Actually h/o

OpenStudy (anonymous):

y(0) = 2sqrt(5)

OpenStudy (anonymous):

\[y(0)=2 \sqrt5\]

OpenStudy (anonymous):

Yeah i was trying to do that but fluttered it up haha

OpenStudy (anonymous):

Ready?

OpenStudy (anonymous):

Yeah drop some multivariable knowledge on me

OpenStudy (anonymous):

\[(y^2+16)dx+y(3-x)dy=0 \implies -y(3-x)dy=(y^2+16)dx \implies \frac{-y dy}{y^2+16}=\frac{dx}{3-x}\] \[-\int\limits \frac{y dy}{y^2+16}=\int\limits \frac{dx}{3-x} \implies -\frac{1}{2}\ln(y^2+16)=\ln(3-x)+C\] Find C. Then solve for Y.

OpenStudy (anonymous):

There isn't any multivar here :P

OpenStudy (anonymous):

Well I certainly was looking at it the wrong way lol. Just out of curiosity can you integrate (y^2+16)dx at all like it is? You can't can you?

OpenStudy (anonymous):

You could. \[\int\limits y^2+16 dx=xy^2+16x+C\]

OpenStudy (anonymous):

Oh, well that's exactly what I wrote down but I figured that wasn't the right way to solve this problem

OpenStudy (anonymous):

Because: \[\frac{\partial}{\partial x}(xy^2+16x+C)=y^2+16\]

OpenStudy (anonymous):

Its not. The way I did is the right way^^ You can't just run a derivative across a sum of differentials like that. Integration is a linear operator meaning: \[\int\limits(a+b)=\int\limits a+\int\limits b \] That does not imply: \[a+b=0 \implies \int\limits a+\int\limits b=\int\limits 0\] You can't just do that.

OpenStudy (anonymous):

got ya

OpenStudy (anonymous):

Thats like going: \[a+b=2 \implies \sin(a)+\sin(b)=\sin(2)\] Which is not true in general.

OpenStudy (anonymous):

So my next question is how am i supposed to solve this with all those logs?

OpenStudy (anonymous):

It's like Ln(3) + (1/2)(ln(y^2+16)) = C

OpenStudy (anonymous):

Let y=2sqrt(5) since you plugged in zero already for x. Then just use ln properties

OpenStudy (anonymous):

That works out SO fluttering nice bro.

OpenStudy (anonymous):

uhm ... idk how exactly ... i'm still getting something like ln(3)= -ln(36) / 2 + C

OpenStudy (anonymous):

Well: \[\ln(3)+\frac{1}{2}\ln(36) \implies \ln(3)+\ln(36^{\frac{1}{2}}) \implies \ln(3)+\ln(6) \implies \ln(18)\]

OpenStudy (anonymous):

Now i have to write that in "standard conic form" wtf kind of conic involves log functions

OpenStudy (anonymous):

Thats not what they mean by conic form.

OpenStudy (anonymous):

Lol... Yeah I realized that after I typed it. Do you remember what conic form is though i dont remember what it's supposed to look like

OpenStudy (anonymous):

Its the same as standard whatever: http://www.uoguelph.ca/~math2080/pdf/Lectures/w6.pdf

OpenStudy (anonymous):

How can the standard form and standard conic form mean the same when in standard form it is like y' = b(t) whatever

OpenStudy (anonymous):

Theres no y' in the form where it has C in it lol

OpenStudy (anonymous):

Hey just to check my algebra could you put that in explicit form and show me?

OpenStudy (anonymous):

−12ln(y2+16)=ln(3−x)+C

OpenStudy (anonymous):

-(1/2)ln(y^2+16) = ln(3-x) + C that is

OpenStudy (anonymous):

You want to solve that for y?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I'm doin something wrong lol. I can't graph it properly. I either get something outrageously small or I get a line

OpenStudy (anonymous):

Lol this is a tad frustrating I have to admit. Theres nothing more that I hate than doing Honeycutt home work... I'd EASILY take Mrs. Wallins home work